Night Time 20140323
Defined a new target position for
PatCams:
20140322_205100
CalUnitSystem light on
PatCams in
DayTime, TN:
20130322_220800
1:20 - Test of
PockelsCells: 100ns opening time counts: 1900, 900, 280
200ns: 2050, 950, 300
TN:
20140323_012459 solar ligth on M1 -> good illumination to measure
FoV diameter
We put diodes to 0.2A and we scan the Pockels cells opening time: 100ns: 9000, 4000, 1600
500ns: 10200, 4800, 2000
1000ns: 11000, 6000, 2600
1500ns: 12800, 7200, 3300
3:16 - Preset acquire authorizing LUCI@RFBG on ACT0284 (about 4deg from M35, EL 60deg), dichroic slided in, the target is offsetted by [6.0; 1.0]arcsec.
Telescope back on M35, EL 63, AZ 258.
3:31 - We try to recenter the LGS within the
FoV
New IM between LM1 and
PatCams:
20140323_033220
We try to recenter also with a periscope loop
Shell RIP TN:
20140323_035507
JCG removes an actuator from the set. M2 restarted.
TN:
20140323_035833 Pupil loop closed
TN:
20140323_040024 Pupil & jitter loop closed
4:31 - We check the LAN flexures going from 45deg EL to Zentih: at each elevation we recenter the LGS with LM1-PatCams loop and we write down the LM1 motors positions:
EL |
A1 |
A2 |
A3 |
| 45 |
14470 |
8626 |
13265 |
| 50 |
14498 |
8649 |
13212 |
| 60 |
14577 |
8687 |
13095 |
| 70 |
14672 |
8710 |
12972 |
| 80 |
14802 |
8736 |
12818 |
| 90 |
14940 |
8739 |
12679 |
On the way back to EL45 we set on LM1 motors the previous values, then we close the LM1-PatCams loops and we write down the final position.
| EL |
A1 |
A2 |
A3 |
| 80 |
14784 |
8741 |
12828 |
| 70 |
14672 |
8728 |
12962 |
| 60 |
14574 |
8695 |
13092 |
| 50 |
14498 |
8655 |
13207 |
| 45 |
14470 |
8637 |
13256 |
During the descending sweep the LGS are always found within the
PatCams FoV using the ascending LM1 positions.
Another actuator dies and it is removed (560)
TN:
20140323_052208 EL45
TN:
20140323_052542 pupil loop closed, good for subapertures definition
TN:
20140323_052650 jitter loop closed
New slope_null acquired:
20140323_052924
We try to keep going the LM1-PatCams loop
We set a new
PupilRecenteringTarget wrt the lab one: YELLOW = [+0.25; -0.25], RED = [+0.5; -0.5].
20140323_054800
We redefine the subapertures wrt the lab ones: BLUE = [0; -1], YELLOW = [+3; -4], RED = [+5; -7]. New
SubaperturesDefinition: 20140323_055200
New
SlopeOffset: 20140323_055556
TN:
20140323_055940 half laser is closed
TN:
20140323_062326 good
LAN focus scan:
| Stage [steps] |
TN |
Notes |
| 0k |
20140323_060444 |
|
| -10k |
20140323_060518 |
|
| -20k |
20140323_060546 |
|
| -30k |
20140323_060643 |
FWHM: 2.2, 1.8 |
| -40k |
20140323_060902 |
|
| -60k |
20140323_061010 |
FWHM: 1.95 |
| -80k |
20140323_061220 |
FWHM: 1.86 |
| -100k |
20140323_061309 |
FWHM: 1.98 |
| -120k |
20140323_061413 |
FWHM: 1.85 |
| -140k |
20140323_061535 |
FWHM: 1.92, 2.1, 1.86, 1.99 |
| -160k |
20140323_061657 |
|
| -180k |
20140323_061852 |
|
Sweep
ToFfirst at big steps then at small ones:
| ToF [us] |
TN |
| 80 |
20140323_062611 |
| 82 |
20140323_062719 |
| 84 |
20140323_062758 |
| 86 |
20140323_062856 |
| 88 |
20140323_062939 |
| 84 |
20140323_063226 |
| 85 |
20140323_063318 |
| 86 |
20140323_063403 |
| 87 |
20140323_063455 |
The best position seems to be
86us
We set incremental modes on M2:
| M2 Z5 [um] |
TN |
SlopeOffset |
| 5 |
20140323_063851 |
20140323_055556 |
| -5 |
20140323_064013 |
20140323_055556 |
| -5 |
20140323_064222 |
20140323_064141 |
| -2.5 |
20140323_064317 |
20140323_064141 |
| -1.25 |
20140323_064408 |
20140323_064141 |
7:00 - Shell RIP, end of games
9:00 - Email to Armando
10:04 - Shell is back, Steve tries to move telescope to EL50, shell RIP again...
11:00 - Test of PI TF with sinusoidal command
| Commanded freq [Hz] |
TN |
Notes |
| 1 |
20140323_114547 |
A=1.35" |
| 2 |
20140323_115229 |
|
| 4 |
20140323_115753 |
wrong |
| 4 |
20140323_120918 |
3.5Hz |
| 8 |
20140323_121157 |
6.5Hz |
| 16 |
20140323_121423 |
12Hz, A=1.3" |
| 32 |
20140323_121620 |
20Hz, A=1.25" |
| 64 |
20140323_121843 |
26Hz, A=1.25" |
Test with a step command.